No relevant resource is found in the selected language.

This site uses cookies. By continuing to browse the site you are agreeing to our use of cookies. Read our privacy policy>

Reminder

To have a better experience, please upgrade your IE browser.

upgrade

Ethernet Link fluctuations in OSN 7500 ASON network, where backbone is through Optix BWS 1600 G

Publication Date:  2012-07-25 Views:  53 Downloads:  0
Issue Description
There were two issues where trails fluctuated regularly, which reported B3_SD alarm at NE 2818 for few seconds, and the services were interrupted at that time. 
At same time, the NE 06 reported B3_SD alarm also.
All the equipments are OSN 7500 in ASON Mesh network and backbone is through Optix BWS 1600G.
ASON GCP V1R5
NE version: V1R8
Alarm Information
B3_SD, HP_UNEQ, AU_AIS
Handling Process
From the SCNP events:
2010-12-3  13:16:47Source:DL_DELH_06XXX_M01_W_L_HDXC001-15-N1SF64-1(SDH-
1)-VC4-23;Sink:DL_DELH_06XXX_M01_W_L_HDXC001-8-N1SLD64-1(SDH-1)-VC4-14;Status: working at protection channel;Switching Status: SF Switching;working channel current status: SF;protection channel current 
The log means that there is a SF on the working channel. When there is a SF on the working channel and protection channel reported B3_SD, the SNCP wont 
switching to the working channel.
At the working channel there is alarm for AU-AIS, it reported SF.
AU_AIS DL_DELH_06XXX_M01_W_L_HDXC001 15-N1SF64-1(SDH-1)-VC4:23 03/12/2010 11:05:34 03/12/2010 14:21:06
 As the B3_SD is not a re-route condition, the ASON re-route didn’t happen also at that time.
  So the B3_SD was generated in ASON network and pass through to the ECI network, then reported at NE 2818. 
There were many bit error & error seconds SDH performance in OSN equipment recently, needed to improve fiber performance to stop fluctuations
Root Cause
One trail information and down time are shown as below:
We check the events of this trail recently and B3_SD reported for several seconds on this trail every day.
According analyze the data, we find the B3_SD and bit error performance reported from NE DL_DELH_06XXX_M01_W_L_HDXC001 to NE GJ_AHMD_04-1X_M01_W_L_HDXC001.
RJ_BWDI_20XXX_B01_W_L_HDXC001 6 N1SF64 1 4 VC4 49 48 HPBBE 15M 2010-12-03 13:30:00+05:30 429 block 
RJ_BWDI_20XXX_B01_W_L_HDXC001 6 N1SF64 1 4 VC4 49 49 HPES     15M 2010-12-03 13:30:00+05:30 53 s 
RJ_BWDI_20XXX_B01_W_L_HDXC001 6 N1SF64 1 4 VC4 49 51 HPFEBBE 15M 2010-12-03 13:30:00+05:30 9057 block 
RJ_BWDI_20XXX_B01_W_L_HDXC001 6 N1SF64 1 4 VC4 49 52 HPFEES 15M 2010-12-03 13:30:00+05:30 573 s 
RJ_BWDI_20XXX_B01_W_L_HDXC001 6 N1SF64 1 4 VC4 49 53 HPFESES 15M 2010-12-03 13:30:00+05:30 34 s 
RJ_BWDI_20XXX_B01_W_L_HDXC001 6 N1SF64 1 4 VC4 49 54 HPUAS 15M 2010-12-03 13:30:00+05:30 14 s 
RJ_BWDI_20XXX_B01_W_L_HDXC001 26 N1SF64 1 4 VC4 49 51 HPFEBBE 15M 2010-12-03 13:30:00+05:30 246 block 
RJ_BWDI_20XXX_B01_W_L_HDXC001 26 N1SF64 1 4 VC4 49 52 HPFEES 15M 2010-12-03 13:30:00+05:30 28 s 
RJ_BWDI_20XXX_B01_W_L_HDXC001 26 N1SF64 1 4 VC4 49 53 HPFESES 15M 2010-12-03 13:30:00+05:30 9 s 
So we know the problem should be at NE GJ_AHMD_04-1X_M01_W_L_HDXC001 and check the performance of the optical port.
GJ_AHMD_04-1X_M01_W_L_HDXC001 4 N1SF64 RS 1 1 RSBBE 15M 2010-12-03 13:30:00+05:30 8576 block PMP_BBE 
GJ_AHMD_04-1X_M01_W_L_HDXC001 4 N1SF64 RS 1 2 RSES 15M 2010-12-03 13:30:00+05:30 132 s PMP_ES 
GJ_AHMD_04-1X_M01_W_L_HDXC001 4 N1SF64 RS 1 6 RSUAS 15M 2010-12-03 13:30:00+05:30 14 s PMP_UAS 
According the performance records list as above, we know there is UAS on the fiber link between 13:15 and 13:30. The optical channel is fluctuating at that time. 
This is the reason that the problem happened.
Suggestions
 None.

END